Tavinona Company issued 1,000; $1,000 bonds for a bond issue of $1,000,000. The bonds were issued for a 10-year term with a coupon rate of 4.5%. a. The current market rate is 5.8%. Calculate the present (market) value of one bond, then calculate the amount of the bond discount.

  • $902.40; $97.60

  • $912.19; $87.81

  • $949.86; $50.14

  • $944.27; $55.73
Answer #1
Assuming semi-annual interest payments:
Amount PV factor Present value
Semi-annual Interest 22.50 15.01596 337.86
Principal 1000 0.56454 564.54
902.40
Present (market) value 902.40
Amount of the bond discount 97.60 =1000-902.40
Option A $902.40; $97.60 is correct
Workings:
Semi-annual Interest 22.50 =1000*4.5%/2
Semi-annual market rate 2.90% =5.8%/2
Number of periods 20 =10*2
PV factor:
Semi-annual Interest 15.015961 =(1-(1.029)^-20)/0.029
Principal 0.564537 =1/1.029^20